OpenStudy (anonymous):

lim (1+logx-x)/(1-2x+x^2) = ? x->1

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1} \frac{ 1+logx - x }{ 1-2x+x^2 }\]

OpenStudy (anonymous):

now expand log x

OpenStudy (anonymous):

log x, expansion? what is it ? can you please tell me

OpenStudy (anonymous):

okk leave it

OpenStudy (anonymous):

u knw l'hopital rule ??

OpenStudy (anonymous):

differentiate numerator and denominator seperatly

OpenStudy (anonymous):

Yes I did that and got -1/2 as the answer. But it is not there in the options given.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1} \frac{ 0+\frac{ 1 }{ x } -1}{ 0-2+2x }\]

OpenStudy (anonymous):

its a silly Q but u type the Q correctly??

OpenStudy (anonymous):

it is log x or ln x

OpenStudy (anonymous):

Its lnx

OpenStudy (anonymous):

Please provide the final answer to it. I got it -1/2.

OpenStudy (anonymous):

its coming -1/2 only

OpenStudy (anonymous):

Ok, with the expansion too I am getting -1/2. So may be its the mistake of the options given for the question. Thank you very much.